# Language Repository for Long Video Understanding

### v1

@@ 5.2.0.3 IntentQA:
IntentQA [42] is based on the same NExT-QA videos (specifically, the ones for temporal and causal reasoning), yet re-purposed focusing on intent-related questions. A new set of 16k multiple-choice questions are provided for 4.3k videos, each with 5 answer choices. These questions are further classified as why?, how? or before/after, based on the appearance of the same keywords in the question. In our evaluation, we focus on zero-shot setting on test set of 2k questions, reporting performance for each category.
@@ 5.2.0.4 NExT-GQA:
NExT-GQA [81] is a visually-grounded VQA dataset, which provides additional 10.5K temporal grounding annotations on top of original NExT-QA [80] data. The test split contains 5.5k question-answer pairs, together with grounding labels that convey the temporal span needed to answer the question. Here, we focus on zero-shot inference similar to [92], and report multiple metrics including Intersection-over-Prediction (IoP) that measures the overlap w.r.t. the predicted window, Intersection-over-Union (IoU) that measures the overlap w.r.t. the union of ground-truth and predicted windows, and Acc@GQA that measures the accuracy of correctly-grounded predictions.
5.3 Main results
Table 2: Results on EgoSchema [51]: We present comparisons with state-of-the-art models on EgoSchema subset (500-videos) and fullset (5000-videos). We focus on the zero-shot setting. LangRepo shows a strong performance at its scale. Models with significantly higher #parameters or video-caption ${}^{\dagger}$ are de-emphasized.
Model Params  Subset (%)  Fullset (%)
finetuned
MC-ViT-L [3] 424M 62.6 44.4
ImageViT [56] 1B 40.8 30.9
ShortViViT [56] 1B 47.9 31.0
LongViViT [56] 1B 56.8 33.3
zero-shot
VIOLET[19] 198M - 19.9
InternVideo [76] 478M - 32.1
FrozenBiLM [86] 890M - 26.9
[…]
mPLUG-Owl [88] 7B - 31.1
LLoVi [92] 7B 50.8 33.5
Vamos [74] 13B - 36.7
[…]
LangRepo (ours) 7B 60.8 38.9
LangRepo (ours) 12B 66.2 41.2
5.3.0.1 EgoSchema:
In Table 2, we present the VQA performance of LangRepo on standard EgoSchema [51] splits, comparing with other state-of-the-art frameworks. Here, we focus on zero-shot evaluation, yet also report finetuned setting (i.e., any downstream-data-specific training) for completeness. We consider Mistral-7B [29] and Mixtral-8$\times$7B [30] as the choice of LLMs in our setup, both with reasonable model scales (7B and 12B active parameters, respectively). We de-emphasize models with significantly-higher #parameters (e.g. 175B GPT-3.5, or 1.8T GPT-4 variants), and the ones that use video-caption pretraining. LangRepo shows significantly-better performance compared to other methods at a similar scale, validating its effectiveness in long-video reasoning. We achieve $+10.1\%$ on fullset compared to mPLUG-Owl [88], $+15.4\%$ on subset and $+7.7\%$ on fullset compared to LLoVi [92] (w/ Mistral-7B [29]), and $+4.5\%$ on fullset compared to Vamos [74] (w/ Llama2-13B [73]).
Table 3: Results on NExT-QA [80]: We compare LangRepo against state-of-the-art methods on NExT-QA validation set, highlighting standard splits: causal, temporal and descriptive. We focus on the zero-shot setting. Our method shows strong performance at its scale. Models with significantly higher #parameters or video-caption ${}^{\dagger}$ are de-emphasized.
 Model Params  Causal (%)  Temporal (%)  Descriptive (%)  All (%)
finetuned
CoVGT [84] 149M 58.8 57.4 69.3 60.0
SeViTFiD [34] 215M - - - 60.6
HiTeA [87] 297M 62.4 58.3 75.6 63.1
MC-ViT-L [3] 424M - - - 65.0
InternVideo [76] 478M 62.5 58.5 75.8 63.2
BLIP-2 [43] 4B 70.1 65.2 80.1 70.1
[…]
LLama-VQA [35] 7B 72.7 69.2 75.8 72.0
Vamos [74] 7B 72.6 69.6 78.0 72.5
zero-shot
VFC [53] 164M 45.4 51.6 64.1 51.5
InternVideo [76] 478M 43.4 48.0 65.1 49.1
[…]
Mistral [29] 7B 51.0 48.1 57.4 51.1
LLoVi [92] 7B 55.6 47.9 63.2 54.3
[…]
LangRepo (ours) 7B 57.8 45.7 61.9 54.6
LangRepo (ours) 12B 64.4 51.4 69.1 60.9
5.3.0.2 NExT-QA:
In Table 3, we report the performance of LangRepo on standard NExT-QA [80] validation splits (Causal, Temporal and Descriptive) and the full validation set. On zero-shot evaluation, our framework outperforms other methods consistently. Compared to smaller models, we gain $+11.8\%$ over InternVideo [76] and $+9.4\%$ over VFC [53]. Compared to models of similar scale, we gain $+9.8\%$ over baseline Mistral-7B [29] and $+6.6\%$ over LLoVi (w/ Mistral-7B) [92]. We de-emphasize LLoVi variant with GPT-4 due to its whopping scale of 1.8T parameters. We also de-emphasize SeViLA [90] which is pretrained with video-captions, whereas we rely on LLaVA-1.5 [48] captions that has not seen any video pretraining. Finally, we observe that LangRepo outperforms competition on semantic splits showing the generalization of our language representation.
Table 4: Results on IntentQA [42]: We compare LangRepo against state-of-the-art methods on IntentQA test set, highlighting standard splits: why?, how? and before/after. We focus on the zero-shot setting. Our method shows strong performance at its scale. Models with significantly higher #parameters or video-caption ${}^{\dagger}$ are de-emphasized.
Model Params  Why? (%)  How? (%)  Before/After (%)  All (%)
finetuned
HQGA [82] 46M 48.2 54.3 41.7 47.7
VGT [83] 511M 51.4 56.0 47.6 51.3
Vamos [74] 7B 69.5 70.2 65.0 68.5
[…]
zero-shot
[…]
Mistral[29] 7B 52.7 55.4 41.5 50.4
LLoVi [92] 7B 57.9 55.4 42.3 53.6
[…]
LangRepo (ours) 7B 56.9 60.2 42.1 53.8
LangRepo (ours) 12B 62.8 62.4 47.8 59.1
5.3.0.3 Intent-QA:
In Table 4, we evaluate our zero-shot framework against other state-of-the-art models on Intent-QA [42] test splits (Why?, How? and Before/After) and the full test set. LangRepo outperform comparable models with similar scale consistently, showing gains of $+8.7\%$ over baseline Mistral-7B [29] and $+5.5\%$ over LLoVi [92] (w/ Mistral-7B). Again, we de-emphasize significantly larger models (LLoVi w/ GPT-4) and models pretrained with video-captions (SeViLA [90]), whereas ours rely on LLaVA-1.5 [48] frame-level captions.
Table 5: Results on NExT-GQA [81]: We compare LangRepo against state-of-the-art methods on NExT-GQA test set. We focus on the zero-shot setting. Our method shows strong performance at its scale. Models with significantly higher #parameters or video-caption ${}^{\dagger}$ are de-emphasized.
 Model Params mIoP IoP@0.5 mIoU IoU@0.5 Acc@GQA
weakly-supervised
IGV [45] 110M 21.4 18.9 14.0 9.6 10.2
Temp[CLIP] [59, 81] 130M 25.7 25.5 12.1 8.9 16.0
FrozenBiLM [86, 81] 1B 24.2 23.7 9.6 6.1 17.5
[…]
zero-shot
Mistral [29] 7B 20.4 20.2 8.7 5.9 9.2
LLoVi [92] 7B 20.7 20.5 8.7 6.0 11.2
[…]
LangRepo (ours) 7B 20.3 20.0 8.7 6.0 11.2
LangRepo (ours) 12B 31.3 28.7 18.5 12.2 17.1
5.3.0.4 NExT-GQA:
In Table 5, we compare the performance of LangRepo with state-of-the-art models on NExT-GQA. Our method achieves a strong performance at its scale, outperforming baseline Mistral-7B [29] by $+7.9\%$ and LLoVi [92] (w/ Mistral-7B) by $+5.9\%$ on Acc@GQA metric. Despite being zero-shot, it is also competitive with weakly-supervised baselines. Here, we de-emphasize significantly-larger models (LLoVi w/ GPT-4) and models pretrained with video-captions (SeViLA [90]).
@@ 5.4 Ablation study
Table 6: Ablations on EgoSchema [51]: We evaluate different design decisions of our framework on EgoSchema 500-video subset for zero-shot video VQA.
LLM   Scale Acc. (%)
Llama2 [73] 13B 43.0
Mistral [29] 7B 50.8
Text encoder Acc. (%)
Sentence-T5-XL [62]
56.4
CLIP-L/14 [59]
57.8
VQA classifier       Acc. (%)
@@
#Iterations  #Chunks  Scales-read  Acc. (%)
1 [2] [1st] 57.0
1 [4] [1st] 60.8
2 [3,2] [2nd] 55.6
2 [3,2] [1st, 2nd] 60.6
3 [4,3,2] [3rd] 58.4
3 [4,3,2] [2nd, 3rd] 59.4
3 [4,3,2] [1st, 2nd, 3rd] 61.2
Model
Acc. (%)
LangRepo (ours)
60.8
+\;\;++ timestamps
60.4
+\;\;++ occurrences
61.4
+\;\;++ timestamps $+$ occurrences
58.2
Streaming setup Acc. (%)
LLoVi [92]
50.8
Chunk-based LLoVi
57.8
LangRepo (ours)
60.8
Model
0.5$\times$
1$\times$
2$\times$
Mistral-7B [29]
49.8
48.8
46.8
LLoVi [92]
57.2
55.4
53.6
LangRepo (ours)
56.4
57.8
56.4
@@ 5.4.0.1 Choice of backbone LLM, text encoder and classifier:
We consider different open-source LLM-backbones in our setup, namely, LLama2 [73] and Mistral [29]. We ablate the choice of LLM within LLoVi [92] framework, as shown in Table LABEL:tab:ablation:llm_choice. We observe that Mistral-7B is significantly better at video reasoning compared to LLama2-13B. Next, we consider different text encoders to embed our text descriptions prior to pruning. Namely, we use either CLIP-L/14 [59] or Sentence-T5-XL [62]. As presented in Table LABEL:tab:ablation:text_encoder. Surprisingly, CLIP outperforms Sentence-T5 that is trained with a sentence-level objective (which is expected to better align with our caption-similarity computation). Finally, we evaluate different classifiers used for close-ended (i.e., multiple-choice question) VQA setups (see Table LABEL:tab:ablation:vqa_classifier). Even though the generative classifier is commonly-used in LLM literature, we observe that the log-likelihood classifier performs better. Such performance is also intuitive as the latter constrains predictions within the given answer choices (hence, no hallucination). Based on this observation, we adopt the log-likelihood classifier in our main experiments.
@@ 5.4.0.2 Repository setup and metadata:
In the formulation of LangRepo we ablate different hyperparameter settings related to the number of repo-updates (#iterations), the number of video chunks in each iteration (#chunks), and multiple temporal-scales considered when reading data in repository. In Table LABEL:tab:ablation:memory_setup, we make two observations: (1) more update iterations with finer chunks (higher #chunks per iteration) can preserve more-useful information, and (2) reading information in multiple temporal-scales is consistently better. Moreover, we consider optional metadata that can help preserve information that may get lost when pruning (e.g. temporal ordering, or repetitive captions), namely, timestamps and #occurrences (i.e., the number of captions grouped within each repo description). We see in Table LABEL:tab:ablation:metadata that #occurrences help weigh each description when summarizing, resulting in better performance. However, timestamps do not provide meaningful improvement in our setup, in the context of EgoSchema VQA.
@@ 5.4.0.3 Input format and length:
We also consider different ways of consuming long video data, either as a whole or as chunks (see Table LABEL:tab:ablation:streaming_setup). Among these options, processing as chunks enables preserving more fine-grained details in LLM outputs. Our repository setup provides further improvement showing its effectiveness over the baseline with the same chunk-based processing. Finally, we re-visit the experiment on how the input length affects the effectiveness of LLMs, presented in Table 1. In Table LABEL:tab:ablation:input_length, we show that LangRepo provide more-stable performance with increasing input lengths, in contrast to baselines, validating how it addresses this shortcoming.
@@ 6.0.0.1 Acknowledgements:
This work was supported by the SUNY Empire Innovation Program (EIP). We thank the Robotics Lab at SBU for helpful discussions, and especially Xiang Li, for managing the lab servers and promptly responding to all the infrastructure-related issues.
A Appendix
A.1 Prompting for VQA
As the evaluation setup, we consider multiple-choice visual question-answering (VQA) on long videos. Given the close-ended answer formulation, we can consider two different classifiers to make the prediction: (1) a Generative classifier, which directly generates the answer choice, or (2) a Log-likelihood classifier, which select the most-probable choice based on the joint-probability of tokens in each answer option given the description and the question. As we discussed in Sec. 4.3, the latter generally performs better, as it is less-prone to hallucinations (i.e., prediction is explicitly constrained to answer choices). However, it is also sensitive to the prompts we use. Hence, we include a discussion on prompting in the following subsections.
A.1.0.1 Generative classifier:
Here, we direcly prompt the LLM to generate the correct answer, conditioned on the descriptions generated by LangRepo, the question and the answer options (inspired by [92]). To make sure that the output can be parsed, we provide additional guiding instructions and any syntax specific to the LLM (Mistral [29]). This also discourages any hallucinations. On all benchmarks, we use the common prompt given below.

### v2

@@ 5.1.0.3 IntentQA:
In Table 5.1.0.2, we evaluate our zero-shot framework against other state-of-the-art models on IntentQA (Li et al., 2023a) test splits (Why?, How? and Before/After) and the full test set. LangRepo outperform comparable models with similar scale consistently, showing gains of $+3.4\%$ over baseline Mistral LLM (Jiang et al., 2023) and $+2.5\%$ over LLoVi (12B) (Zhang et al., 2023a). Again, we de-emphasize significantly larger models and those pretrained with video-captions.
Table 5: Results on NExT-GQA (Xiao et al., 2023a): We compare LangRepo against state-of-the-art methods on NExT-GQA test set. We focus on the zero-shot setting. Our method shows strong performance at its scale.
{tabu}
lccccccc  Model Video Pretrain Params mIoP IoP@0.5 mIoU IoU@0.5 Acc@GQA
weakly-supervised
IGV (Li et al., 2022) ✓ 110M 21.4 18.9 14.0 9.6 10.2
Temp[CLIP] (Radford et al., 2021; Xiao et al., 2023a) ✓ 130M 25.7 25.5 12.1 8.9 16.0
FrozenBiLM (Yang et al., 2022; Xiao et al., 2023a) ✓ 1B 24.2 23.7 9.6 6.1 17.5
SeViLA (Yu et al., 2024) ✓ 4B 29.5 22.9 21.7 13.8 16.6
zero-shot (with proprietary LLMs)
MoReVQA (Min et al., 2024) ✗ 340B 37.8 37.6 19.7 15.4 39.6
LLoVi (Zhang et al., 2023a) ✗ 1.8T 37.3 36.9 20.0 15.3 24.3
zero-shot (with open-source LLMs)
Mistral (Jiang et al., 2023) ✗ 7B 20.4 20.2 8.7 5.9 9.2
LLoVi (Zhang et al., 2023a) ✗ 7B 20.7 20.5 8.7 6.0 11.2
LLoVi (Zhang et al., 2023a) ✗ 12B 31.4 28.8 18.4 12.0 16.2
LangRepo (ours) ✗ 7B 20.3 20.0 8.7 6.0 11.2
LangRepo (ours) ✗ 12B 31.3 28.7 18.5 12.2 17.1
@@ 5.1.0.4 NExT-GQA:
In Table 5.1.0.3, we compare the performance of LangRepo with state-of-the-art models on NExT-GQA (Xiao et al., 2023a). We follow the same grounding setup as in (Zhang et al., 2023a). Our method achieves a strong performance at its scale, outperforming baseline Mistral LLM (Jiang et al., 2023) by $+2.0\%$ and LLoVi (12B) (Zhang et al., 2023a) by $+0.9\%$ on Acc@GQA metric. Despite being zero-shot, it is also competitive with weakly-supervised baselines. Here, we de-emphasize significantly-larger models and those pretrained with video-captions.
@@ 5.2 Ablation study
Table 6: Ablating design decisions on EgoSchema (Mangalam et al., 2024): We evaluate different design decisions of our framework on EgoSchema 500-video subset for zero-shot video VQA.
{tabu}
lcc LLM Scale Acc.
Llama2 (Touvron et al.) 13B 43.0
Mistral (Jiang et al.) 7B 50.8
{tabu}
lc Text encoder Acc.
Sentence-T5-XL (Reimers & Gurevych) 56.4
CLIP-L/14 (Radford et al.) 57.8
{tabu}
L3.4cmc VQA classifier Acc.
@@
{tabu}
C1cmC1.5cmcc #Iter #Ch Read Acc.
1 [2] 1 57.0
1 [4] 1 60.8
3 [4,3,2] 1 58.4
3 [4,3,2] 2 59.4
3 [4,3,2] 3 61.2
{tabu}
L3.3cmc Model Acc.
LangRepo (ours) 60.8
+\;\;++ tstmp 60.4
+\;\;++ occ 61.4
+\;\;++ tstmp $+$ occ 58.2
{tabu}
lcccc Model Params Latency per video (s)
q/v = 1 q/v = 2 q/v = 5
LLoVi (Zhang et al.) 7B 22.11 44.34 108.75
LangRepo 7B 30.98 37.46 56.90
LLoVi (Zhang et al.) 12B 50.06 99.84 249.95
LangRepo 12B 85.09 94.90 124.33
{tabu}
L3.5cmc Captions Acc.
BLIP-2 (Li et al.) 55.4
LLaVA-1.5 (Liu et al.) 58.4
LaViLa (Zhao et al.) 60.8
Oracle 69.2
{tabu}
L4.2cmc Streaming setup Acc.
LLoVi (Zhang et al.) 50.8
Chunk-based LLoVi 57.8
LangRepo (ours) 60.8
{tabu}
lC0.55cmC0.55cmC0.55cm Model 0.5$\times$ 1$\times$ 2$\times$
Mistral (Jiang et al.) 49.8 48.8 46.8
LLoVi (Zhang et al.) 57.2 55.4 53.6
LangRepo 56.4 57.8 56.4
@@ 5.2.0.1 Choice of backbone LLM, text encoder and classifier:
We ablate the choice of LLM-backbones within the framework in Zhang et al. (2023a) in Table LABEL:tab:ablation:llm_choice. We observe that Mistral-7B (Jiang et al., 2023) is significantly better at video reasoning compared to LLama2-13B (Touvron et al., 2023). Next, we consider different text encoders to embed our text descriptions prior to pruning, such as CLIP-L/14 (Radford et al., 2021) or Sentence-T5-XL (Reimers & Gurevych, 2019) in Table LABEL:tab:ablation:text_encoder. Surprisingly, CLIP outperforms Sentence-T5 that is trained with a sentence-level objective (which is expected to better align with our caption-similarity computation). Finally, we evaluate different classifiers used for close-ended (i.e., multiple-choice question) VQA setups (see Table LABEL:tab:ablation:vqa_classifier). Despite commonly-used in LLM literature, generative classifier performs worse than log-likelihood classifier. Such performance is also intuitive as the latter constrains predictions within the given answer choices (hence, less hallucination). More discussion on this is in supplementary.
@@ 5.2.0.2 Repository setup and metadata:
In the formulation of LangRepo we ablate different hyperparameter settings related to the number of repo-updates (#iterations), the number of video chunks in each iteration (#chunks), and multiple temporal-scales considered when reading data in repository. In Table LABEL:tab:ablation:memory_setup, we make two observations: (1) more update iterations with finer chunks (higher #chunks per iteration) can preserve more-useful information, and (2) reading information in multiple temporal-scales is consistently better. Moreover, we consider optional metadata to help preserve information that may get lost when pruning (e.g. temporal ordering, or repetitive captions), namely, timestamps and #occurrences (i.e., the number of captions grouped within each repo description). We see in Table LABEL:tab:ablation:metadata that #occurrences help weigh each description when summarizing, resulting in better performance. However, timestamps do not provide meaningful improvement in our setup, in the context of EgoSchema VQA.
5.2.0.3 Efficiency in a multi-query setup:
We also ablate the efficiency of our concise representation in Table LABEL:tab:ablation:efficiency. LangRepo can be initially expensive, as it requires multiple write-read operations (yet, each processing smaller context-lengths). However, once repository is created, it can be re-used more-efficiently in a setup with multiple-queries for a given video (i.e., the initial cost will be amortized). This is especially relevant in practical scenarios, where users may have multiple queries correponding to a given video.
5.2.0.4 Captioner quality:
In Table LABEL:tab:ablation:captioner, we evaluate the quality of captions consumed by LangRepo. By default, we use short-clip captions from LaViLa (Zhao et al., 2023c), which outperform frame-level captions (BLIP-2 (Li et al., 2023b), LLaVA-1.5 (Liu et al., 2023)). Oracle captions from Ego4D show the performance upper-bound.
@@ 5.2.0.5 Input format and length:
We consider different ways of consuming long video data, either as a whole or as chunks (see Table LABEL:tab:ablation:streaming_setup). Among these options, processing as chunks enables preserving more fine-grained details in LLM outputs. Our repository setup provides further improvement showing its effectiveness over the baseline with the same chunk-based processing. Finally, we re-visit the experiment on how the input length affects the effectiveness of LLMs, presented in Table 3. In Table LABEL:tab:ablation:input_length, we show that LangRepo provide more-stable performance with increasing input lengths, in contrast to baselines.
@@ 6.0.0.1 Acknowledgements:
This work was supported by Electronics and Telecommunications Research Institute (ETRI) grants funded by the Korean government [24ZR1100, A Study of Hyper-Connected Thinking Internet Technology by autonomous connecting, controlling and evolving ways]. This work was also supported by the Institute of Information & communications Technology Planning & Evaluation (IITP) grant funded by the Korea government (MSIT) [No. RS-2024-00336738, Development of Complex Task Planning Technologies for Autonomous Agents, 100%]. The authors thank the Robotics Lab at SBU for helpful discussions, and especially Xiang Li, for managing the lab servers and promptly responding to all the infrastructure-related issues.
Reproducibility Statement
We use open-source LLMs (w/ publicly-available code and pretrained-weights) in all our experiments. By relying on LLMs with reasonable-scale (i.e., not proprietary, paid LLMs), we make our work more-accessible. As all our experiments are done in zero-shot settings, we do not update any pretrained weights. All our evaluations are conducted on publicly-available standard long-video benchmarks. We detail all required steps, and provide prompts to reproduce the proposed contributions. Finally, we release our code together with the paper to support further research.
Appendix A Appendix
A.1 Design decisions
A.1.0.1 Similarity-based pruning:
A.1.0.2 Processing videos as chunks:
Our decision to consume longer videos as chunks is motivated by prior work (Wu et al., 2022; Ryoo et al., 2023). It allows us to not lose short-term details, while also keeping track of long-term dependencies via multi-scale processing. Additionally, although not explored in the scope of this paper, such a setup integrates well with temporally-fine-grained prediction tasks, where an LLM needs to make multiple predictions over time.
A.1.0.3 Choice of metadata:
A.1.0.4 All-textual repository:
Instead of being a latent representation (Wu et al., 2022; Ryoo et al., 2023; Balažević et al., 2024), our LangRepo is all-textual. This promotes interpretability for human observers, while also being a more-natural form of structure for LLM-based processing. Additionally, our implementation can be formulated to be zero-shot, without requiring any training or finetuning.
A.1.0.5 Classifier for close-ended VQA:
The standard multiple-choice question-answering setup considers a generative classifier. Meaning, an LLM is prompted to generate the correct answer option among multiple-choices, directly as next-token prediction. Another approach used in NLP literature is log-likelihood based classification (see Cloze prompting in (Robinson et al., 2023)). Here, the LLM is prompted separately for each of the multiple choices with a template such as “Question: Answer-option”. The choice that maximises the log-likelihood of predicted tokens (i.e., tokens corresponding to Answer-option) is selected as the correct answer. This is a more-natural setup for close-ended VQA since it avoids hallucination. Among these classifiers, we find the latter to be better-performing. Yet, it is more-sensitive to the prompt template. We direct the reader to supplementary A.2 for more details.
A.2 Prompting for VQA
As the evaluation setup, we consider multiple-choice visual question-answering (VQA) on long videos. Given the close-ended answer formulation, we can consider two different classifiers to make the prediction: (1) a Generative classifier, which directly generates the answer choice, or (2) a Log-likelihood classifier, which select the most-probable choice based on the joint-probability of tokens in each answer option given the description and the question. As we discussed in Sec. A.1, the latter generally performs better, as it is less-prone to hallucinations (i.e., prediction is explicitly constrained to answer choices). However, it is also sensitive to the prompts we use. Hence, we include a discussion on prompting in the following subsections.
A.2.0.1 Generative classifier:
Here, we direcly prompt the LLM to generate the correct answer, conditioned on the descriptions generated by LangRepo, the question and the answer options (inspired by (Zhang et al., 2023a)). To make sure that the output can be parsed, we provide additional guiding instructions and any syntax specific to the LLM (Mistral (Jiang et al., 2023)). This also discourages any hallucinations. On all benchmarks, we use the common prompt given below.
{adjustwidth}